An inheritance of $15,000 is to be split and invested at 6% and 9% interest to produce an average 8% interest rate. How much money is to be invested at each rate?
Answer by richwmiller(17219)

0.08*15000=$1200 interest earned on $15000 at 8%
Total amount of money invested: $15000
x+y=15000,
Total yearly interest for the two accounts is: $1200
0.06*x+0.09*y=1200
x=15000-y
Substitute for x
0.06*(15000-y)+0.09*y=1200
Multiply out
900-0.06*y+0.09*y=1200
Combine like terms.
0.03*y=300
Isolate y
y=$10000.00 at 9%
x=15000-y
Calculate x
x=$ 5000.00 at 6%
Check
0.06*5000+0.09*10000=1200
300+900=1200
1200=1200
If this statement is TRUE and neither x nor y is negative then all is well
code int1